OpenStudy (anonymous):

Find the particular solution of the differential equation f''(x)=6(x-1) whose graph passes through (2,1) and is tangent to the line 3x-y-5=0 at that point

OpenStudy (anonymous):

i know how to find the integral but im wondernig about the first C value

OpenStudy (calculusfunctions):

Would you like me to teach you @rjeffries96 ?

OpenStudy (anonymous):

@calculusfunctions SHE REALLY NEEDS HELP AND SHE HASTO GET OFF AT 630 sorry for the caps

OpenStudy (anonymous):

im just confused about how to find the first c by using the tangent line portion of the problem (thanks mikala1)

OpenStudy (calculusfunctions):

Alright but @mikala1 you know me well enough now to know that I'm not simply going to just give her the answer? BTW Hi!

OpenStudy (anonymous):

haha hi :) so am i right when i say that the first derivative and the equation given as the tangent line will be the same line?

OpenStudy (calculusfunctions):

@rjeffries96 no problem can you write down the answer you have thus far so that I can see that you've done it correctly?

OpenStudy (anonymous):

well right now i have that the integral of 6x-6 = 3x^2 -6x +C

OpenStudy (anonymous):

i need help finding the C value, then i know to take the 2nd integral and sub in the point given to yield f(x)

OpenStudy (calculusfunctions):

Now you know that f '(x) is the slope. You also know that since the tangent to this curve is 3x - y - 5 = 0 The slope of this tangent equals f '(x).

OpenStudy (anonymous):

Ok, that makes sense. But then how would I find C? would i set the two equations equal (I tried and it wasnt much use) or do i substitute in the values x=2 and y=1?

OpenStudy (calculusfunctions):

No the point (2, 1) is for f(x). All you know at the moment is that f '(x) = 3. Before I proceed further, are you sure you wrote the entire question, as is? Double check please?

OpenStudy (anonymous):

Yes, that is the entire question

OpenStudy (anonymous):

how did you find that f'(x)=3?

OpenStudy (calculusfunctions):

Because the slope of the given tangent line is 3.

OpenStudy (anonymous):

oh ok (oops i shoulve seen that)

OpenStudy (anonymous):

its 6 30 sorry to interupt

OpenStudy (anonymous):

im really sorry but i have to be at school in 10 minutes, i can just ask my teacher. thanks anyways

OpenStudy (anonymous):

well did he help you in any way

OpenStudy (calculusfunctions):

OK when you're back on next time, let me know and I'll show you if you still don't understand.

OpenStudy (anonymous):

thanks for helping him/her lol she/him told me to tell u but she was late ololo

OpenStudy (calculusfunctions):

It's Ok @mikala1 I'll break the rules just this once and post the solution for her.

OpenStudy (anonymous):

i think she got off but when ever she gets on

OpenStudy (calculusfunctions):

f ''(x) = 6x - 6 f '(x) = 3x² - 6x + C We are told that 3x - y - 5 = 0 is tangent to the curve at the point (2, 1). This implies that both the curve and the tangent line pass through this point. Hence the first derivative (slope) is 3 at x = 2. Thus if f '(x) = 3x² - 6x + C then 3 = 3(2)² - 6(2) + C 3 = 12 - 12 + C C = 3 Thus f '(x) = 3x² - 6x + 3 Now f(x) = x³ - 3x² + 3x + K Now f(x) passes through (2, 1). Thus 1 = (2)³ - 3(2)² + 3(2) + K 1 = 8 - 12 + 6 + K K = -1 ∴ f(x) = x³ - 3x² + 3x - 1

OpenStudy (calculusfunctions):

@mikala1 if for some reason you're able to find her right now, please ask her to check it out.

OpenStudy (anonymous):

ok she left for school so she will be on when she gets home she said but i think it is great and easy to understand

OpenStudy (calculusfunctions):

Thank you @mikala1 I have to sign out now but before I do, do you need anything?

OpenStudy (anonymous):

no thank you byyyy

OpenStudy (calculusfunctions):

BYE!!!